# Real-time Bayesian inference at extreme scale:

### v1

@@ VIII Implications
We enable real-time probabilistic tsunami forecasts for near-field events such as a future magnitude 8–9 CSZ megathrust earthquake, where conventional systems may fail due to delayed or inaccurate source characterization. Our digital twin framework resolves complex rupture dynamics and tsunami wavefields at unprecedented speed and accuracy, enabling early predictions of inundation and evacuation zones within seconds, critical for life-saving response where warning time is otherwise measured in minutes or less. Notably, the real-time inversion itself requires only moderate computational resources. Moreover, if only surface wave heights at selected locations are of interest, the forecasting step reduces to a precomputed, small, dense matrix–vector product—enabling deployment entirely without any HPC infrastructure.
While we demonstrate the potential for real-time forecasting, we remain limited by the sparsity of offshore sensors currently available in the CSZ; however, emerging technologies such as distributed acoustic sensing will improve observational coverage for resolving near-field tsunami source characteristics [Schmidt2023, becerril2024towards]. Furthermore, expanding to fully-coupled acoustic–elastic simulations [krenz2021palu] allows us to employ our framework to invert for fault slip, and forward propagate seismic waves to compute—in real time—maps of the intensity of ground motion in populated regions. These shake maps provide critical information for early responders and post-earthquake recovery.
More generally, autonomous dynamical systems arise in many different settings beyond geophysical inversion. Our Bayesian inversion-based digital twin framework is thus more broadly applicable to such problems as acoustic, electromagnetic, and elastic inverse scattering; source inversion for transport of atmospheric or subsurface hazardous agents; satellite inference of emissions; and treaty verification, among others.
@@ Acknowledgment
The authors thank Pythagoras Watson and Livermore Computing for assistance with El Capitan runs, Amanda Dufek and Muaaz Awan for assistance with Perlmutter runs, John Cazes for assistance with Frontera runs, Jonatan Glehman for assistance in generating earthquake simulation output, and Eric Dunham for insightful discussions about this work.
[…]

### v2

@@ VIII Implications
We enable real-time probabilistic tsunami forecasts for near-field events such as a future magnitude 8–9 CSZ megathrust earthquake, where conventional systems may fail due to delayed or inaccurate source characterization. Our performance-portable digital twin framework resolves complex rupture dynamics and tsunami wavefields at unprecedented speed and accuracy, enabling early predictions of inundation and evacuation zones within seconds, critical for life-saving response where warning time is otherwise measured in minutes. Notably, the real-time inversion itself requires only moderate computational resources. Moreover, if only surface wave heights at selected locations are of interest, the forecasting step reduces to a precomputed, small, dense matrix–vector product—enabling deployment entirely without any HPC infrastructure.
While we demonstrate the potential for real-time forecasting, we remain limited by the sparsity of offshore sensors currently available in the CSZ; however, emerging technologies such as distributed acoustic sensing will improve observational coverage for resolving near-field tsunami source characteristics [33, 34]. Furthermore, expanding to fully-coupled acoustic–elastic simulations [35] allows us to employ our framework to invert for fault slip, and forward propagate seismic waves to compute—in real time—maps of the intensity of ground motion in populated regions. These shake maps provide critical information for early responders and post-earthquake recovery.
More generally, autonomous dynamical systems arise in many different settings beyond geophysical inversion. Our Bayesian inversion-based digital twin framework is thus more broadly applicable to acoustic, electromagnetic, and elastic inverse scattering; source inversion for transport of atmospheric or subsurface hazardous agents; satellite inference of emissions; and treaty verification, among others.
@@ Acknowledgment
The authors thank Pythagoras Watson and Livermore Computing for assistance with El Capitan runs, Daniel Ganellari and Sebastian Keller for assistance with Alps runs, Amanda Dufek and Muaaz Awan for assistance with Perlmutter runs, John Cazes for assistance with Frontera runs, Jonatan Glehman for assistance in generating earthquake simulation output, and Eric Dunham for insightful discussions about this work.